# The sPHENIX Micromegas Outer Tracker

### v1

@@ Abstract
The sPHENIX Time Projection Chamber Outer Tracker (TPOT) is a Micromegas based detector. It is a part of the sPHENIX experiment that aims to facilitate the calibration of the Time Projection Chamber, in particular the correction of the time-averaged and beam-induced distortion corrections. This paper describes the detector mission, setup, construction, installation, commissioning and performance during the first year of sPHENIX data taking.
@@ 1 Introduction
The sPHENIX detector PHENIX:2015siv ; Belmont:2023fau is an experiment located at the Relativistic Heavy Ion Collider (RHIC) HARRISON2003235 in Brookhaven National Laboratory (BNL). It focuses on measuring jets as well as open and hidden heavy flavor production in heavy ion collisions and to study the properties of the Quark Gluon Plasma created in gold on gold (${\rm Au+Au}$) collisions at a center of mass energy per nucleon-nucleon collision ${\sqrt{s_{\rm NN}}}=200$ GeV. It will also collect data using proton-proton (${p+p}$) and possibly proton-gold (${p+{\rm Au}}$) collisions to serve as a reference to the ${\rm Au+Au}$ data and to study the cold nuclear matter that constitutes the initial state of such collisions.
@@
TPOT is a late addition to the sPHENIX apparatus. It is installed on the outside of the TPC and greatly facilitates the calibration of the TPC. In particular, it allows to make the maximum use of tracks for measuring in near real time the distortions of the electron drift in the TPC during data taking. It consists of 16 Micromesh Gaseous detectors (Micromegas) GIOMATARIS199629 grouped two-by-two to provide an additional space point on the outside of the TPC in a fraction of the TPC acceptance. Figure 2 presents a schematic view of how the TPC can be calibrated using all of the other tracking detectors. A reference track (in red) is built using the clusters measured in the MVTX, INTT and TPOT (filled red point). It is interpolated inside the TPC volume and the expected cluster positions are inferred (open red circles). For each TPC layer, the expected position is compared to the measured clusters in the TPC and the difference between the two is used to calibrate the TPC. This calibration is extrapolated to the full acceptance of the TPC.
@@ 2.1 Requirements
Performance: because TPOT is not used as a measurement device critical for the sPHENIX physics program but rather as a calibration device for the TPC, the requirements on the detector in terms of detection efficiency and spatial resolution are somewhat loose. A minimum of 90% detection efficiency is required, together with a spatial resolution of a few 100 $\mu$m in both azimuthal and longitudinal directions. The positioning of the detector itself is not critical, and an accuracy of 1 mm or less is required. This is provided that the real position (alignment) of the detector can be measured a posteriori with a precision of 100 $\mu$m or less, significantly smaller than the expected spatial resolution of the detector.
@@
Cost and schedule: the decision to construct and install TPOT as a part of the sPHENIX apparatus was taken late in the sPHENIX design process, namely about eighteen months before the beginning of sPHENIX operations, and based on design studies started approximately six months before that. This compressed schedule, together with the total budget allocated to the project put strong constraints on the amount of prototyping that could be performed before starting the construction of the detector, as well as the number of modules that could be built. In particular, it was decided early on that TPOT would only cover a fraction of the TPC acceptance (Section 2.2). Extrapolation methods would then be developed offline to extrapolate the calibrations provided by TPOT to the rest of the TPC acceptance.
@@
Magnetic field: TPOT is installed inside sPHENIX superconducting magnet, which delivers an approximately uniform longitudinal magnetic field of intensity 1.4 T. The magnetic field directly impacts TPOT operation and in particular the choice of gas that is used to operate the Micromegas chambers. In addition, its presence prevents the use of any magnetic component for both the detector and the mechanical support structure.
@@
Particle rates and discharges: because of the requirements above in terms of performances, cost, schedule and available space, it was realized early on that the Micromegas technology GIOMATARIS199629 is a good candidate for building TPOT (section 2.3). However such detectors are known to be susceptible to electrostatic discharges when the number of electrons created in the avalanche process responsible for the signal amplification exceeds the Raether limit raether1964electron of $2\times 10^{7}$ electrons. Based on simulations of the expected particle rates in TPOT, the mean deposited energy of these particles, and for a typical amplification gain of $6\times 10^{3}$, the rate of such electrostatic discharge was estimated to be about 100 Hz. To prevent these discharges from generating a prohibitively high dead-time in the detector or even damaging the readout electronics, they must be quenched, in the TPOT case by adding a resistive layer on top of the readout plane.
@@ 2.2 Detector segmentation
TPOT is made up of eight modules each consisting of two Micromegas chambers stacked radially, one to measure the coordinates along the longitudinal direction ($z$) and the second along the azimuthal direction ($\phi$). They are referred to as the $z$ and $\phi$ views, respectively. The dimension of each chamber’s active area is 256 mm along $\phi$ and 512 mm along $z$. The signal on the $\phi$ views is collected on 256 strips oriented along $z$ and with a pitch of 1 mm. The signal on the $z$ views is collected on 256 strips oriented along $\phi$ and with a pitch of 2 mm. Four of the eight TPOT modules cover the full $z$ extent of the bottom-most sector of the TPC (out of twelve sectors total). The other four modules are located along the neighbor two sectors (two modules each). With this configuration, the eight TPOT modules cover approximately 8% of the TPC acceptance.
@@ 2.3.1 Principle
Figure 3 shows a schematic view of a bulk resistive Micromegas chamber. Each chamber consists of a drift electrode, a micromesh, a resistive layer and a readout layer arranged in parallel top to bottom. In order to guide electrons through the drift (or conversion) gap, a negative electric field is applied through the layers. The drift electrode is operated at negative voltage. The micromesh is located 3 mm away from the drift electrode and connected to the electrical ground. The resistive layer is located 128 $\mu$m away from the micromesh and is operated at a positive voltage. The resistive layer is segmented into four groups to be connected to the High Voltage (HV) independently. This segmentation allows for mitigation of a HV short in one part of the chamber while allowing the other 3/4th of the chamber to remain active. The readout layer is located immediately below the resistive layer and consists of 256 parallel strips per chamber. The strips read the signal collected on the resistive layer through capacitive coupling and are connected to the Front-End Electronics (FEE) boards.
@@
To optimize the fabrication process, the chambers have no components soldered on the readout board. The thickness of the readout Printed Circuit Board (PCB) is 1.6 mm to match the SAMTEC MEC8 edge connector standard. One side of the readout PCB is designed to form 4 MEC8 edge connectors to connect the 256 strips to the front-end electronics.
@@
The chambers are circulated with a 95/5% mixture of Ar/iC4H10 which provides good gain and stability for TPOT operation. The distribution of the gas to the inside of the chamber is carried out by the 3D-printed frame that also holds the drift electrode and an O-ring seal for gas tightness. This 5 mm thick frame cannot accommodate standard L-shape gas connectors. 3 mm diameter polyurethane tubes glued directly inside the frame are used instead. The tubes are capped with filters to ensure cleanliness inside the chambers at all times (Figure 5). Each chamber is operated at a gas flow rate ranging from $50$ to 100 ccm and slightly above atmospheric pressure.
@@ 2.3.2 The readout electrode
Figure 6: Layout of the readout PCB for the $\phi$ (left) and the $z$ (left) view. Shown is the layer in which readout strips are routed to 4 MEC8 connectors located at the bottom.
2.3.3 The resistive layer
The resistive layer has several purposes: protection against electromagnetic discharges, protection of the readout electronics, spread of the charge, and rate limitation. Here the main concern is caused by the many highly ionizing particles created in heavy ion collisions. Early estimates of the charge deposited in the drift volume of the chamber and assuming a ${\rm Au+Au}$ collision rate of 50 kHz gave a 100 Hz rate for crossing the Raether limit raether1964electron and likely causing discharges. Hence, the resistive layer was deemed necessary to protect the SAMPA electronics. At 50 kHz collision rate, the resistive layer must evacuate the electric charges quickly, which forbids the use of Diamond-Like Carbon (DLC) ROBERTSON1992185 as a resistive layer. With low resistivity comes large cluster sizes. To prevent this, the layer is segmented in the same direction as the readout strips. A number of pitch values and patterns have been tested for the resistive layer segmentation, illustrated in Figure 7. The one with the smallest pitch (500 $\mu$m) was found to lead to the best (smallest) spatial resolution and was adopted for both views.
Figure 7: Examples of resistive layer patterns tested on a 2 mm straight strip readout pattern. From left to right : (i) 2 mm pitch and 1 mm straight strip; (ii) 1 mm pitch and 600 $\mu$m straight strip; (iii) 500 $\mu$m pitch and 290 $\mu$m straight strip and (iv) 2 mm zigzag strips.
@@ 2.3.4 Fabrication
The fabrication process of a typical bulk Micromegas chamber is described in Giomataris_2006 . For TPOT additional steps are needed to include the resistive layer (Section 2.3.3). The resistive layer itself is fabricated using a thin layer of isolating material (Kapton) on which the strip pattern is printed using serigraphy. The readout PCB (Section 2.3.2) and the resistive layer are then glued and pressed together. The gluing and pressing is performed at CERN’s Micro-Pattern Gaseous Detector (MPGD) laboratory. The resulting assembly, several layers of photoresistive film (Vacrel) of the appropriate thickness and the woven wire micromesh are then laminated together at high temperature. The photoresistive material is subsequently etched using photolitography to produce the pillars that keep the micromesh at the proper distance (128 $\mu$m) from the resistive layer (Figure 3). These steps are performed in the MPGD laboratory at the CEA research center of Saclay. In parallel, the carbon fiber supported drift electrode is mounted on the 3D-printed frame which defines the chamber active volume and provides gas circulation. Both are assembled to the Micromegas bulk, using a set of screws and an O-ring seal between the bulk and the printed frame to ensure gas tightness.
@@ 2.3.5 Conditioning
At Saclay, the Micromegas bulk-assembly (readout plane, resistive layer and micromesh, Figure 3) is first conditioned in air and without a drift electrode, to a voltage up to 800 V. HV boards (Section 2.3, Figure 4) are tested up to 1000 V for approximately twelve hours. The drift electrodes are conditioned independently up to 1500 V. This conditioning is repeated once the chamber is assembled and the gas tightness is verified, using the nominal 95/5% Ar/iC4H10 gas mixture, up to a value of 450 V for the resistive layer, and 500 V for the drift electrode. This corresponds to a detector amplification gain of approximately $5\times 10^{4}$, a value significantly larger than needed for routine data taking in sPHENIX (Sec. 4.3.2). Gas tightness is checked at 2 l/h gas flow and a leak rate tolerance of 0.02 l/h.
@@ 2.4 Mechanical structure
Figure 8 shows a schematic view of a TPOT module, with electronic boards and cooling plates mounted. Each module consists of two Micromegas chambers (one $\phi$ and one $z$ view) mounted back to back, two FEE boards (one per chamber) and two water cooled aluminum plates.
Figure 8: Schematics of a fully assembled TPOT module.
@@ 2.5.1 Readout system architecture
Figure 10 shows the block diagram of the readout system for TPOT. The signal from TPOT is fed into the FEE which amplifies, shapes and digitizes it into data. Digital data is sent to the back-end electronics which consists of a Data Aggregation Module (DAM) installed inside the Event Buffering and Data Compressor (EBDC). The DAM is a specialized version of the FELIX board developed for the ATLAS experiment at the Large Hadron Collider (LHC) Ryu_2017 . The EBDC is a high-end rack-mounted commodity server. Compressed data from the EBDC is sent to Buffer Boxes, which are high capacity, high input-output disks, and serve as a buffer before it is transferred to the Scientific Data and Computing Center (SDCC) at BNL and recorded on permanent storage (tapes).
@@ 2.5.2 The 32-channel SAMPA ASIC
The ASIC used for the FEE is called SAMPA v5 Barboza_2016 ; 8772086 . Primarily designed for the ALICE TPC and Muon Chamber upgrades Adolfsson_2021 ; 9507849 , the ASIC was modified for the sPHENIX TPC requirement, the main modification consisting of including a shaping time of 80 ns in place of 320 ns. A description and block diagram of the ASIC is found in Barboza_2016 .
For TPOT, the SAMPA operation parameters are: (i) 160 ns shaping time, which is the longest available value, chosen to increase the collection of the ion tail in the Micromegas; (ii) 20 mV/fC gain, which is the smallest available gain, chosen to minimize the fraction of signal that would saturate the ADC range and (iii) 50 MHz sampling rate.

### v3

@@ Abstract
The sPHENIX Time Projection Chamber Outer Tracker (TPOT) is a Micromegas based detector. It is a part of the sPHENIX experiment that aims to facilitate the calibration of the Time Projection Chamber, in particular the correction of the time-averaged and beam-induced distortions of the electron drift. This paper describes the detector mission, setup, construction, installation, commissioning and performance during the first year of sPHENIX data taking.
@@ 1 Introduction
The sPHENIX detector [1, 2] is an experiment located at the Relativistic Heavy Ion Collider (RHIC) [3] in Brookhaven National Laboratory (BNL). It focuses on measuring jets as well as open and hidden heavy flavor production in heavy ion collisions and to study the properties of the Quark Gluon Plasma created in gold on gold (${\rm Au+Au}$) collisions at a center of mass energy per nucleon-nucleon collision ${\sqrt{s_{\rm NN}}}=200$ GeV. It will also collect data using proton-proton (${p+p}$) and possibly proton-gold (${p+{\rm Au}}$) collisions to serve as a reference to the ${\rm Au+Au}$ data and to study the cold nuclear matter that constitutes the initial state of such collisions.
@@
TPOT is a late addition to the sPHENIX apparatus. It is installed on the outside of the TPC and greatly facilitates the calibration of the TPC. In particular, it allows to make the maximum use of tracks for measuring in near real time the distortions of the electron drift in the TPC during data taking. It consists of 16 Micromesh Gaseous detectors (Micromegas) [4] grouped two-by-two to provide an additional space point on the outside of the TPC in a fraction of the TPC acceptance. Figure 2 presents a schematic view of how the TPC can be calibrated using all of the other tracking detectors. A reference track (in red) is built using the clusters measured in the MVTX, INTT and TPOT (filled red point). It is interpolated inside the TPC volume and the expected cluster positions are inferred (open red circles). For each TPC layer, the expected position is compared to the measured clusters in the TPC and the difference between the two is used to calibrate the TPC. This calibration is then extrapolated to the full acceptance of the TPC.
@@ 2.1 Requirements
Performance: because TPOT is not used as a measurement device critical for the sPHENIX physics program but rather as a calibration device for the TPC, the requirements on the detector in terms of detection efficiency and spatial resolution are somewhat loose. A minimum of 90% detection efficiency is required, together with a spatial resolution of a few 100 $\mu$m in both azimuthal and longitudinal directions. The requirement on the detector spatial resolution is driven by the amount of time it takes to achieve an accuracy of order 10 $\mu$m for reconstructing the distortions of the electron drift in the TPC with sufficient granularity (a few square centimeters). With 500 $\mu$m resolution, a trigger rate of 15 kHz and an expected mean number of particles per Micromegas detector and ${\rm Au+Au}$ collision around eight, this integration time is estimated to be less than a minute, much smaller than the typical time over which the distortions of the electron drift in the TPC is expected to vary significantly, beyond event by event fluctuations. The positioning of the detector itself is not critical, and an accuracy of 1 mm or less is required. This is provided that the real position (alignment) of the detector can be measured a posteriori with a precision of 100 $\mu$m or less, significantly smaller than the expected spatial resolution of the detector.
@@
Cost and schedule: the decision to construct and install TPOT as a part of the sPHENIX apparatus was taken late in the sPHENIX design process, about eighteen months before the beginning of sPHENIX operations, and based on design studies started approximately six months before that. This compressed schedule, together with the total budget allocated to the project put strong constraints on the amount of prototyping that could be performed before starting the construction of the detector, as well as the number of modules that could be built. In particular, it was decided early on that TPOT would only cover a fraction of the TPC acceptance (Section 2.2). Extrapolation methods would then be developed offline to extrapolate the calibrations provided by TPOT to the rest of the TPC acceptance.
@@
Magnetic field: TPOT is installed inside sPHENIX superconducting magnet, which delivers an approximately uniform longitudinal magnetic field of intensity 1.4 T. The magnetic field directly impacts TPOT operation, in particular the choice of gas that is used to operate the Micromegas chambers and the electric field in the detector’s drift gap (Section 2.3.1). In addition, its presence prevents the use of any magnetic component for both the detector and the mechanical support structure.
@@
Particle rates and discharges: because of the requirements above in terms of performance, cost, schedule and available space, it was realized early on that the Micromegas technology [4] is a good candidate for building TPOT (section 2.3). However such detectors are known to be susceptible to electrostatic discharges when the number of electrons created in the avalanche process responsible for the signal amplification exceeds the Raether limit [5] of $2\times 10^{7}$ electrons. Based on simulations of the expected particle rates in TPOT, the mean deposited energy of these particles, and for a typical amplification gain of $6\times 10^{3}$, the rate of such electrostatic discharge was estimated to be about 100 Hz. To prevent these discharges from generating a prohibitively high dead-time in the detector or even damaging the readout electronics, they must be quenched, in the TPOT case by adding a resistive layer on top of the readout plane.
@@ 2.2 Detector segmentation
TPOT is made up of eight modules each consisting of two Micromegas chambers stacked radially, one to measure the coordinate along the longitudinal direction ($z$) that follows the beam axis and the second along the azimuthal direction ($\phi$), as defined in polar coordinates around $z$. They are referred to as the $z$ and $\phi$ views, respectively. The dimension of each chamber’s active area is 256 mm along $\phi$ and 512 mm along $z$. The signal on the $\phi$ views is collected on 256 strips oriented along $z$ and with a pitch of 1 mm. The signal on the $z$ views is collected on 256 strips oriented along $\phi$ and with a pitch of 2 mm. Four of the eight TPOT modules cover the full $z$ extent of the bottom-most sector of the TPC (out of twelve sectors total), and approximately one half of its azimuthal extent. The other four modules are located along the neighbor two sectors (two modules each). With this configuration, the eight TPOT modules cover approximately 8% of the TPC acceptance.
@@ 2.3.1 Principle
Figure 3 shows a schematic view of a bulk resistive Micromegas chamber, and Figure 4 a schematic view of a fully assembled TPOT module (out of eight modules total), with electronic boards and cooling plates mounted.
@@
Figure 4: Schematics of a fully assembled TPOT module.
Each chamber consists of a drift electrode, a micromesh, a resistive layer and a readout layer arranged in parallel top to bottom. In order to guide electrons through the drift (or conversion) gap, a negative electric field is applied through the layers. The drift electrode is operated at negative voltage. The micromesh is located 3 mm away from the drift electrode and connected to the electrical ground. The resistive layer is located 128 $\mu$m away from the micromesh and is operated at a positive voltage. It is segmented into four groups to be connected to the High Voltage (HV) independently. This segmentation allows for mitigation of a HV short in one part of the chamber while allowing the other 3/4th of the chamber to remain active. The readout layer is located immediately below the resistive layer and consists of 256 parallel strips per chamber. The strips read the signal collected on the resistive layer through capacitive coupling and are connected to the Front-End Electronics (FEE) boards.
To optimize the fabrication process, the chambers have no components soldered on the readout board. The thickness of the readout Printed Circuit Board (PCB) is 1.6 mm to match the SAMTEC MEC8 edge connector standard. One side of the readout PCB is designed to form 4 MEC8 edge connectors to connect the 256 strips to the front-end electronics (Figure 7).
@@
The chambers are circulated with a 95/5% mixture of Ar/iC4H10 which provides good gain and stability for TPOT operation. In addition, this mixture has a small Lorentz angle, estimated to be less than 15°$\degree$ in sPHENIX conditions, and defined as the angle between the drift direction of the electrons and the direction of the electric field in presence of a longitudinal magnetic field. It also allows to operate at a large drift electric field (about 1 kV/cm for TPOT). The distribution of the gas to the inside of the chamber is carried out by the 3D-printed frame that also holds the drift electrode and an O-ring seal for gas tightness. This 5 mm thick frame cannot accommodate standard L-shape gas connectors. 3 mm diameter polyurethane tubes glued directly inside the frame are used instead. The tubes are capped with filters to ensure cleanliness inside the chambers at all times (Figure 6). Each chamber is operated at a gas flow rate ranging from $50$ to 100 cc/min (cubic centimeter per minute) and slightly above atmospheric pressure. The material used for the 3D-printed frame is a high quality plastic, Accura ClearVue Free, SL 7870.
@@ 2.3.2 The readout electrode
The readout electrode is a multi-layer PCB of thickness 1.6 mm and size 316$\times$542 mm2, with an active area of 256$\times$512 mm2. The top layer corresponds to the chamber readout strips. For $\phi$ views, there are 256 straight strips of length $512$ mm along the long dimension of the PCB ($z$) and with a pitch of 1 mm and an interstrip distance of 100 $\mu$m. For the $z$ views, there are 256 strips of length $256$ mm along the short dimension of the PCB ($\phi$) and with a pitch of 2 mm and an interstrip distance of 500 $\mu$m. A zigzag pattern is chosen for the $z$ view strips in order to increase the electric charge sharing between neighbor strips and thus the cluster size in the chamber, despite the large pitch. The other layers of the PCB are used for routing the readout strips to 4 MEC8 connectors of 64 (active) + 6 (ground) pads each, located on the side of the PCB (Figure 7). A ground layer is also added to the PCB layout to provide appropriate electromagnetic shielding to the chamber.
Figure 7: Layout of the readout PCB for the $\phi$ (left) and the $z$ (right) view. Shown are the strips that collect the signal, the four MEC8 connectors located at the bottom, as well as the location where the HV board from Fig. 5 connects to the detector. The actual pitch of the detector strips (1 mm for the $\phi$ view and 2 mm for the $z$) is not properly represented on the figures, due to their finite resolution.
2.3.3 The resistive layer
@@
Figure 8: Examples of resistive layer patterns tested on a 2 mm straight strip readout pattern. From left to right : (i) 2 mm pitch and 1 mm straight strip; (ii) 1 mm pitch and 600 $\mu$m straight strip; (iii) 700 $\mu$m pitch and 290 $\mu$m straight strip and (iv) 2 mm zigzag strips.
@@ 2.3.4 Fabrication
The fabrication process of a typical bulk Micromegas chamber is described in [6]. For TPOT additional steps are needed to include the resistive layer (Section 2.3.3). The final stack consists of the PCB, a kapton with the printed resisitive ink, two layers of 64 $\mu$m photoresistive film (Pyralux), the mesh, a third layer of photoresistive film of thickness 64 $\mu$m and the drift plate. The resistive layer itself is fabricated using a thin layer of isolating material (Kapton) on which the strip pattern is printed using ink screening technique. The readout PCB (Section 2.3.2) and the resistive layer are then glued (EM-370 pre-preg from Elite Material Co., Ltd.) and pressed together. The gluing and pressing is performed at CERN’s Micro-Pattern Gaseous Detector (MPGD) laboratory. The resulting assembly is laminated twice at high temperature. The photoresistive material is subsequently etched using photolitography to produce the pillars that keep the micromesh at the proper distance (128 $\mu$m) from the resistive layer (Figure 3). The pillars are cylindrical, with a diameter of 400 $\mu$m, and located every 4 mm (3 mm) along (perpendicular to) $z$. The woven wire micromesh is then laminated on top of the stack and the third photoresistive layer is etched so to only keep the frame of the mesh and hold it into position. These steps are performed in the MPGD laboratory at the CEA research center of Saclay. A 45-18 stainless steel mesh from BOPP SD is used, with 18 $\mu$m diameter wires and 45 $\mu$m aperture, corresponding to 400 lines per inch. In parallel, the carbon fiber supported drift electrode is mounted on the 3D-printed frame which defines the active volume of the chamber and provides gas circulation. The drift electrode is made of a polyimide film laminated with a copper foil, glued onto a carbon plate for sturdiness. A second polyimide film laminated with a copper foil is glued to the top of the carbon plate, for grounding. The drift electrode mounted on the 3D-printed frame is then assembled to the Micromegas bulk, using a set of screws and an O-ring seal between the bulk and the printed frame to ensure gas tightness.
@@ 2.3.5 Conditioning
At Saclay, the Micromegas bulk-assembly (readout plane, resistive layer and micromesh, Figure 3) is first conditioned in air and without a drift electrode, to a voltage up to 800 V. HV boards (Section 2.3, Figure 5) are tested up to 1000 V for approximately twelve hours. The drift electrodes are conditioned independently up to 1500 V. This conditioning is repeated once the chamber is assembled and the gas tightness is verified, using the nominal 95/5% Ar/iC4H10 gas mixture, up to a value of 450 V for the resistive layer, and 500 V for the drift electrode. This corresponds to a detector amplification gain of approximately $5\times 10^{4}$, a value significantly larger than needed for routine data taking in sPHENIX (Sec. 4.3.2). Gas tightness is checked at 2 l/h gas flow and a leak rate tolerance of 0.05 l/h.
@@ 2.5.1 Readout system architecture
Figure 10 shows the block diagram of the readout system for TPOT. The signal from TPOT is fed into the FEE which amplifies, shapes and digitizes it into data. Digital data is sent to the back-end electronics which consists of a Data Aggregation Module (DAM) installed inside the Event Buffering and Data Compressor (EBDC). The DAM is a specialized version of the FELIX board developed for the ATLAS experiment at the Large Hadron Collider (LHC) [7]. The EBDC is a high-end rack-mounted commodity server. Compressed data from the EBDC is sent to Buffer Boxes, which are high capacity, high input-output disks, and serve as a buffer before it is transferred to the Scientific Data and Computing Center (SDCC) at BNL and recorded on permanent storage (tapes).
@@ 2.5.2 The 32-channel SAMPA ASIC
The ASIC used for the FEE is called SAMPA v5 [8, 9]. Primarily designed for the ALICE TPC and Muon Chamber upgrades [10, 11], the ASIC was modified for the sPHENIX TPC requirement, the main modification consisting of including a shaping time of 80 ns in place of 320 ns. A description and block diagram of the ASIC is found in [8].
For TPOT, the SAMPA operation parameters are: (i) 160 ns shaping time, which is the longest available value, chosen to increase the collection of the ion tail in the Micromegas; (ii) 20 mV/fC gain, which is the smallest available gain, chosen to minimize the fraction of signal that would saturate the ADC range and (iii) 20 MHz sampling rate.